use the order of operations to simplify the expression
(-3)^2-4^2/6 divided by 3*4-2

Answer by stanbon(57410)

(-3)^2-4^2/6 divided by 3*4-2
= (9-16/6)/(12-2)
= (9-8/3)/10
= 19/3 /10
= 19/30
Cheers
Stan H.